OpenStudy (anonymous):

The price of a gallon of milk was $2.65. The price rose y dollars after the last hurricane. Then the price dropped $0.15 and later rose again by $0.05. Which expression represents the current price of milk?

OpenStudy (anonymous):

Price = 2.65 Price = 2.65 + y Price = 2.65 + y - 0.15 Price = 2.65 + y - 0.15 + 0.05 These are the steps illustrated in the equation. Solving the Final Price: Price = y +2.55

OpenStudy (anonymous):

thnkx